OpenStudy (anonymous):

If 0.25 + x = y and y/x = 0.2, then y =

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (anonymous):

i wanted to use substitution

OpenStudy (anonymous):

and i got -.05

OpenStudy (nuttyliaczar):

I'll give you the next step, see if you can take it from here: .25 + x = y .25/x = y/x

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

u sure thats how u do it?

OpenStudy (nuttyliaczar):

Oh woops forgot the +1

OpenStudy (nuttyliaczar):

.25/x + 1 = y/x